OpenStudy (anonymous):

The measure of an angles supplement is 44 less than the angle. find the measure of the angle and its supplement

OpenStudy (anonymous):

x +(x-44) = 180 2x -44 = 180 2x = 224 x =112 Supplement = 112-44 = 68

OpenStudy (anonymous):

thank you so very much!! :)